Over 2 years, how much more does $2000 in a savings account with an APR of 5.2% compounded quarterly earn in interest than the same amount in a savings account with an APR of 4.8% compounded monthly?
A. $8.31
B. $16.61
C. $4.15
D. $33.22

Answer by rfer(16322)

A=2000(1.013)^2*4
A=2000*1.10885
A=2217.71
-----------------
A=2000(1.004)^2*12
A=2000*1.1005483
A=2201.10
-------------------
2217.71-2201.10=$16.61